Mr. Kim is 25 years older than his daughter. In ten years Mr. Kim will be twice as old as his daughter will be. How old is each now? I know that Mr. Kim is 40 years old and the daughter is 15, but I do not know how to show my work. I do not know how to find the equation.
Answer by Earlsdon(6103)

Let K = Mr Kim's present age and D = his daughter's present age.
From the problem you can write:
1) K = D+25 "Mr Kim is 25 years older than his daughter"
2) K+10 = 2(D+10) "In 10 year, Mr Kim will be twice as old as his daughter will be."
Substitute the K from equation 1) into equation 2) and solve for D.
2b) (D+25)+10 = 2(D+10) Simplify.
2c) D+35 = 2D+20 Subtract D from both sides.
2d) 35 = D+20 Finally, subtract 20 from both sides.
2c) 15 = D or D = 15
Mr Kim's daughter's age is 15 years.
Mr Kim's age is D+25
15+25 = 40 years.